OpenStudy (anonymous):

A light house uses a parabolic reflector that is 1 m in diameter how deep should the reflectpr be if the light source is placed halfwayy between the vertex and the plane of the rim?

OpenStudy (mayankdevnani):

write the equation y=a *x^2 for x= 0.5 y = 0.25 a and y focus ( light source )= 0.125a but in a a parabola y =ax^2+bx+c y focus =(4ac-b^2 +1)/4a so in this case = 1/4a so 1/4a= 0.125a and a^2=2 so a = sqrt2 and depth = 0.35.36 m http://in.answers.yahoo.com/question/index?qid=20101119033031AAmzuSC

OpenStudy (mayankdevnani):

@ali110

OpenStudy (anonymous):

the answer here in the book is y^2 = 36x. i dont understand how to get this

OpenStudy (mayankdevnani):

just round it.

OpenStudy (anonymous):

Answer: 9 inches from the vertex; how?

OpenStudy (mayankdevnani):

36x(approx.)

OpenStudy (anonymous):

can you draw the sketch? :)

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

a large reflecting telescope has a parabolic mirror that is 4 meters in diameter